A rectangular lot is to be enclosed with 500 feet of fencing. The area of the lot is 14,400 square feet. Find the dimensions of the rectangular lot.
Answer by TimothyLamb(4379)

2L + 2w = 500
Lw = 14400
w = 14400/L
---
2L + 2w = 500
2L + 2(14400/L) = 500
2L + 28800/L - 500 = 0
2LL - 500L + 28800 = 0
LL - 250L + 14400 = 0
---
the above quadratic equation is in standard form, with a=1, b=-250 and c=14400
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -250 14400
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the quadratic has two real roots at:
---
L = 160
L = 90
---
the larger root makes better sense as a length rather than a width:
---
L = 160
w = 14400/L
w = 14400/160
w = 90
---
answer:
L = 160 ft
w = 90 ft
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---